OpenStudy (anonymous):

Fayed tipped the pizza delivery driver $3 for pizzas that cost $24. What percent did he tip the driver? 3% 8% 12.5% 21%

OpenStudy (campbell_st):

just use amount of tip/ total cost * 100 = 3/24 * 100 =

OpenStudy (anonymous):

@campbell_st so its 8%

OpenStudy (campbell_st):

nope the calculation is \[3 \div 24 \times 100 = \]

OpenStudy (anonymous):

@campbell_st I got 800 but I thought it was 8 lol